(1 point) Find the solution of Y – 6y +9y = 144 91 with y(0) = 2 and ý (0) = 3. y =
Answer #1

Che y - 67 + gy = 144 Ose y(s) = 2, y(o)=3. Now its auxiliary equation in m? - 6m +9 (m - 3)2 -0 ms 3,3 its complementy soNow from y(0)=2 2 = 4 +4 - C, = -2 4 from y(0)= 3 3=36,+ C₂ + 36 5) C2 = -27 Now the solution is 90+) = (2-427+)234 +487 ryt